OpenStudy (anonymous):

A brick is dropped from a high scaffold. (a) What is its velocity after 4.0s? (b) How far does the brick fall during this time?

OpenStudy (anonymous):

So we know by the word "dropped" its being dropped from rest. so our v initial or vo is zero we are going to use the equation's v(t)=v0+gt and y(t)=yo vot+.5gt^2 a) so vo=0 the object has fallen for 4 seconds vo+gt 0+9.8(4)= 39.2m/s b) Now we used y(t) equation yo we'll set as zero to make things easier and again vo is zero y(t)=0+0+.5(9.8)(4)^2= 78.4m so a) at t=4 the velocity is 39.2m/s downwards and b) at t=4 the brick has fallen 78.4 meters.